Fisk Corporation is trying to improve its inventory control system and has installed an online computer at its retail stores. Fisk anticipates sales of 63,000 units per year, an ordering cost of $4 per order, and carrying costs of $1.40 per unit.

a. What is the economic ordering quantity?



b. How many orders will be placed during the year?



c. What will the average inventory be?
Answer #1

Annual Demand = D = 63000/year
Ordering Cost = Co = $4
Holding Cost = Cc = $1.4

(a) Economic Order Quantity Q* = √(2DCo/Cc) = √(2*63000*4*/1.4) = 600

(b) Number of Orders = D/Q* = 63000/600 = 105

(c) Average Inventory = Q*/2 = = 600/2 = 300